OpenStudy (fullmetal):

cos theta /1 - sin theta = 1 + cos theta + sin theta / 1 + cos theta - sin theta

Directrix (directrix):

Is this an equation to solve or perhaps an identity to prove?

OpenStudy (fullmetal):

prove

Directrix (directrix):

It is not clear which terms are in the numerator and which are in a denominator and which are part of neither. You can click on the blue DRAW button and write the equation by hand or use grouping symbols to clear up the confusion.

Directrix (directrix):

Waiting for a reply. @fullmetal

OpenStudy (anonymous):

\[\frac{\cos (\theta)} {1 - \sin (\theta)} =\frac{ 1 + \cos (\theta) + \sin (\theta)}{ 1 + \cos (\theta) - \sin (\theta)}\]

OpenStudy (fullmetal):

\[\cos \theta/ 1- \sin \theta = 1 + \cos \theta + \sin \theta / 1 + \cos \theta - \sin\]

Directrix (directrix):

@fullmetal It is not clear which terms are in the numerator and which are in a denominator and which are part of neither. You can click on the blue DRAW button and write the equation by hand or use grouping symbols to clear up the confusion. Example: 2 - 5 / 2 - 5  ≠ (2 - 5) / (2 - 5)